# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 2.2 Twisted traces
$q^{k}S\bigg{(}\frac{-v^{k}}{qv-q^{-1}v^{-1}}F(v^{-1})+\frac{v^{-k}}{q^{-1}v-qv^{-1}}F(v)\bigg{)}=q^{-\frac{m}{2}}S\bigg{(}F(qv)\frac{v^{-k-m}}{v-v^{-1}}-F(qv^{-1})\frac{v^{k+m}}{v-v^{-1}}\bigg{)}$ (2.4)
Moreover, since $H_{0},H_{1}$ and $v+v^{-1}$ generate $\mathcal{A}$, any $S$ satisfying (2.4) gives a trace $T$ by $T(\sum w^{i}f_{i})=S(f_{0})$.
Now we check that the integral from the statement of the Proposition satisfies the equation (2.4):
$q^{k}\bigg{(}$ $\int_{\varepsilon S^{1}}\bigg{(}\frac{-z^{k}}{qz-q^{-1}z^{-1}}F(z^{-1})+\frac{z^{-k}}{q^{-1}z-qz^{-1}}F(z)\bigg{)}\omega(z)\frac{dz}{z}\big{)}=$ (2.5)
$q^{-\frac{m}{2}}$ $\big{(}\int_{\varepsilon S^{1}}\bigg{(}F(qz)\frac{z^{-k-m}}{z-z^{-1}}-F(qz^{-1})\frac{z^{k+m}}{z-z^{-1}}\bigg{)}\omega(z)\frac{dz}{z}\bigg{)}.$
@@
$\int_{\varepsilon S^{1}}\bigg{(}F(qz)$ $\frac{z^{-k-m}}{z-z^{-1}}-F(qz^{-1})\frac{z^{k+m}}{z-z^{-1}}\bigg{)}\omega(z)\frac{dz}{z}=$ (2.6)
$\int\limits_{q^{-1}\varepsilon S^{1}}F(qz)$ $\frac{z^{-k-m}}{z-z^{-1}}\omega(z)\frac{dz}{z}-\int\limits_{q\varepsilon S^{1}}F(qz^{-1})\frac{z^{k+m}}{z-z^{-1}}\omega(z)\frac{dz}{z}=$
$\int_{\varepsilon S^{1}}F(z)$ $\frac{q^{k+m}z^{-k-m}}{q^{-1}z-qz^{-1}}\omega(q^{-1}z)\frac{dz}{z}-\int_{\varepsilon S^{1}}F(z^{-1})\frac{q^{k+m}z^{k+m}}{qz-q^{-1}z^{-1}}\omega(qz)\frac{dz}{z}=$
$q^{k+\frac{m}{2}}\bigg{(}\int_{\varepsilon S^{1}}F(z)$ $\frac{z^{-k}}{q^{-1}z-qz^{-1}}\omega(z)\frac{dz}{z}-\int_{\varepsilon S^{1}}F(z^{-1})\frac{z^{k}}{qz-q^{-1}z^{-1}}\omega(z)\frac{dz}{z}\bigg{)}.$
@@
This shows that every function $\omega$ satisfying $\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z)$, $\omega(z)=\omega(z^{-1})$ gives a trace. The space of such functions has dimension $1+\frac{m}{2}$. Indeed, let $\omega(z)=\omega(e^{2\pi iz})$ and $q=e^{2\pi i\tau}$. Then $\omega(z+1)=\omega(z)$, $\omega(-z)=\omega(z)$ and $\omega(z+\tau)=e^{-\pi im(\tau+2z)}$. Any such function can be obtained as $C\prod_{j=1}^{\frac{m}{2}}\vartheta(z-\alpha_{j})\vartheta(z+\alpha_{j})$, where $\vartheta$ is Jacobi theta function and $\alpha_{j}$ and $C$ are any numbers.
Now we prove that all maps $S$ satifying (2.4) can be obtained in this way. It is enough to prove that the dimension of the space of possible $S$ is at most $\frac{m}{2}+1$.
Taking $F$ regular in (2.4) we see that $S(\frac{1}{(v\pm q)(v^{-1}\pm q)})$ is defined by the values of $S$ on regular elements. Suppose that $F$ has a pole at $q^{k}$ for some $k\neq 1$. Then the left-hand side expression has a pole at ±$\pm q^{\pm 1}$ and $q^{\pm k}$ (no pole if $k=0$) and the right-hand side has a pole at $q^{\pm(k-1)}$. It follows by induction that $S(\frac{1}{(v\pm q^{k})(v^{-1}\pm q^{k})})$ is defined by the values of $S$ on regular elements.
Take the same $a\in\mathcal{A}$ as before and $b=-w+w^{-1}$. Writing $T(ba)=T(ag(b))$ in terms of $S$ we get
$S(F(v)+F(v^{-1}))=q^{-\tfrac{m}{2}}S(F(qv)v^{-m}+F(qv^{-1})v^{m}).$
For $F(v)=v^{k}$ we get
$S(v^{k}+v^{-k})=q^{k-\tfrac{m}{2}}S(v^{k-m}+v^{m-k}).$
It follows that $S$ is uniquely defined by $S(1),S(v+v^{-1}),\cdots,S(v^{\frac{m}{2}}+v^{-\frac{m}{2}}).$ Therefore the dimension of possible $S$ is at most $1+\frac{m}{2}$, as claimed.
2.3 Positive traces
Proposition 2.3.1.
Let $T$ be a positive trace. Let $w$ be a function from Proposition 2.2.2. Then $\omega(1)=\omega(-1)=0$.
@@ 2.3 Positive traces
$S(F(z))=\int_{S^{1}}F(z)\omega(z)\frac{dz}{z}+\sum_{i>0}c_{i,\pm}\operatorname{Res}_{\pm q^{i}}F,$
where $c_{i,\pm}$ are nonzero multiples of $\omega(\pm 1)$, so that for at least one of the choices of sign all $c_{i,\pm}$ are nonzero. Without loss of generality assume that $c_{i}=c_{i,+}$ are nonzero.
Let
$a=\big{(}w^{2}F(v)-w^{-2}vF(v^{-1})\big{)}(v-1)^{-1},$
where $F(v)$ is regular. This is an element of $\mathcal{A}$. Then
$\rho(a)=q^{m}\big{(}-w^{2}v^{1-m}\overline{F}(v)+w^{-2}v^{m}\overline{F}(v^{-1})\big{)}(v-1)^{-1}$
@@
$[w^{0}]a\rho(a)=q^{m}\big{(}w^{2}F(v)(v-1)^{-1}w^{-2}v^{m}\overline{F}(v^{-1})(v-1)^{-1}+w^{-2}vF(v^{-1})(v-1)^{-1}w^{2}\overline{F}(v)v^{1-m}(v-1)^{-1}\big{)}.$
@@
$w^{2}F(v)(v-1)^{-1}w^{-2}v^{m}\overline{F}(v^{-1})(v-1)^{-1}$ $+w^{-2}vF(v^{-1})(v-1)^{-1}w^{2}v^{1-m}\overline{F}(v)(v-1)^{-1}=$
$F(q^{2}v)\overline{F}(v^{-1})(q^{2}v-1)^{-1}v^{m}(v-1)^{-1}$ $+F(q^{2}v^{-1})\overline{F}(v)q^{-2}v(q^{-2}v-1)^{-1}v^{1-m}(v-1)^{-1}=$
$=G(v)$ $+G(v^{-1}),$
@@
$G(v)=\overline{F}(v^{-1})F(q^{2}v)\cdot\frac{v^{m}}{(q^{2}v-1)(v-1)}$
@@
$G(v^{-1})=\overline{F}(v)F(q^{2}v^{-1})\cdot\frac{q^{-2}v^{2-m}}{(q^{-2}v-1)(v-1)}.$
@@
$T(a\rho(a))=S([w^{0}]a\rho(a))=S(G(v)+G(v^{-1}))=\int_{\varepsilon S^{1}}\big{(}G(z)+G(z^{-1})\big{)}\omega(z)\frac{dz}{z}.$ (2.7)
@@
$\overline{F}(1)\frac{F(q^{2})}{q^{2}-1}.$
The only pole of $G(z^{-1})$ between 0 and $qS^{1}$ is $v=q^{2}$, and the residue is
$\overline{F}(q^{2})\cdot\frac{F(1)q^{-2}q^{2(2-m)}}{(q^{2}-1)q^{-2}}=F(1)\overline{F}(q^{2})\frac{q^{2(2-m)}}{q^{2}-1}$
Therefore
$\int_{\varepsilon S^{1}}\big{(}G(z)+G(z^{-1})\big{)}\omega(z)\frac{dz}{z}=$ $c_{0}\overline{F}(1)\frac{F(q^{2})}{q^{2}-1}+c_{1}F(1)\overline{F}(q^{2})\frac{q^{2(2-m)}}{q^{2}-1}+$ (2.8)
$\int_{q^{-1}S^{1}}G(z)\omega(z)\frac{dz}{z}+\int_{qS^{1}}G(z^{-1})\omega(z)\frac{dz}{z}$
We rewrite the integrals:
$\int_{q^{-1}S^{1}}G(z)\omega(z)\frac{dz}{z}=$ $\int_{S^{1}}G(q^{-1}z)\omega(q^{-1}z)\frac{dz}{z}=$
$\int_{S^{1}}\overline{F}(qz^{-1})F(qz)\frac{q^{-m}v^{m}}{(qv-1)(q^{-1}v-1)}\omega(q^{-1}z)\frac{dz}{z}.$
$\int_{qS^{1}}G(z^{-1})\omega(z)\frac{dz}{z}=$ $\int_{S^{1}}G(q^{-1}z^{-1})\omega(qz)\frac{dz}{z}=$
$\int_{S^{1}}\overline{F}(qz)F(qz^{-1})\cdot\frac{q^{-m}v^{z-m}}{(v-1)(q^{2}v-1)}\cdot\omega(qz)\frac{dz}{z}$
Now take $F(z)=q^{-N}z^{N}-a$ for some constant $a$ and natural number $N$. For large $N$ the leading term of
$c_{0}\overline{F}(1)\frac{F(q^{2})}{q^{2}-1}+c_{1}F(1)\overline{F}(q^{2})\frac{q^{2(2-m)}}{q^{2}-1}$
is $C_{0}aq^{-N}+C_{1}\overline{a}q^{-N}$, where $C_{0},C_{1}$ are constant numbers. For some choice of $a$ this is a negative number with absolute value growing as $q^{-N}$.
On the other hand, for $z\in S^{1}$ polynomials $\overline{F}(qz^{-1})$, $F(qz)$, $\overline{F}(qz)$, $F(qz^{-1})$ have absolute value at most $1+\lvert a\rvert$. Hence the integrals in the right-hand side of (2.8) are bounded when $N$ tends to infinity. On the other hand, the first two terms in the right-hand side of (2.8) tend to minus infinity. Comparing (2.7) and (2.8) we see that for large enough $N$ we have $T(a\rho(a))<0$, a contradiction with positivity.
In particular, for $m=4$ we get that the function $\omega$ is unique up to a constant. Indeed, consider $\omega$ such that $\omega(z)=\omega(e^{2\pi iz})$. Then $\omega(z)=\vartheta(z-\frac{\tau}{2})\vartheta(z+\frac{\tau}{2})\vartheta(z-\tfrac{1}{2}-\frac{\tau}{2})\vartheta(z+\frac{1}{2}+\frac{\tau}{2})$ up to a constant.
Proposition 2.3.2.
Let $T$ be a positive trace, $a=\sum w^{2i+1}f_{2i+1}(v)$. Then $T(a)=0$.
@@
Using Proposition 2.2.1 it is enough to prove that $T(a)=0$ for elements $a$ of the form $a=wF(v)-wF(v^{-1})$. For these elements we have $T(a)=C_{+}F(\sqrt{q})+C_{-}F(-\sqrt{q})$. Assume that $C_{+}$ is nonzero.
Let $a=wF(v)-w^{-1}F(v^{-1})$ for some Laurent polynomial $F$, $b=a+1$. Then
$\rho(b)=q^{-\frac{m}{4}}\big{(}-wv^{\frac{m}{2}}\overline{F}(v)+w^{-1}v^{-\frac{m}{2}}\overline{F}(v^{-1})\big{)}+1.$
Note that $T(b\rho(b))$ depends only on the $w^{0}$ and $w^{1}$ coefficients of $b\rho(b)$. They can be computed as follows:
$[w^{0}]b\rho(b)=$ $q^{-\frac{m}{4}}\big{(}F(qv)v^{-\frac{m}{2}}\overline{F}(v^{-1})+F(qv^{-1})v^{\frac{m}{2}}\overline{F}(v)\big{)}+1,$
$[w^{1}]b\rho(b)=$ $F(v)-v^{\frac{m}{2}}\overline{F}(v).$
Let $G=F(\sqrt{q}v)$, $I(G)=\int_{S^{1}}\lvert G(z)\rvert^{2}\frac{dz}{z}$. We have $S([w^{0}]b\rho(b))\leq cI(G)$ for some $c>0$. On the other hand, trace for the odd powers of $w$ is a linear combination of $G(1)$, $G(-1)$, $\overline{G}(1)$ and $\overline{G}(-1)$; and $G(1)$ has nonzero coefficient. Let $G_{N}(z)=a(\frac{z+1}{2})^{N}$. Then $G_{N}(1)=a$, $G_{N}(-1)=0$. On the other hand, $G_{N}(z)$ is a sequence of functions such that $\lvert G_{N}(z)\rvert\leq 1$ tends to zero for almost all $z$ as $N$ tends to infinity. It follows that $I(G_{N}(z))$ tends to zero as $N$ tends to infinity. We get that $T(b\rho(b))$ tends to a linear combination of $a$ and $\overline{a}$. We can choose $a$ so that this linear combination does not belong to $\mathbb{R}_{\geq 0}$. We get a contradiction with positivity of $T$. ∎
Let $T$ be a positive trace. Combining Propositions 2.3.1 and 2.3.2 it follows that for $a=\sum w^{i}f_{i}$ we have $T(a)=\int f_{0}(z)\omega(z)\frac{dz}{z}$ and $\omega(1)=\omega(-1)=1$.
Proposition 2.3.3.
Let $T$ be a trace on $\mathcal{A}$ of the form $T(a)=\int f_{0}(z)\omega(z)\frac{dz}{z}$, where $w$ is a holomorphic function on $\mathbb{C}^{*}$ such that $\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z)$, $\omega(z)=\omega(z^{-1})$ and $\omega(1)=\omega(-1)=0$. Then $T$ is positive if and only if $\omega(z)$ and $z^{-\frac{m}{2}}\omega(\sqrt{q^{-1}}z)$ are nonnegative on $S^{1}$.
@@
$\rho(a)=\sum q^{-\frac{k^{2}m}{4}}w^{-k}v^{-\frac{km}{2}}\overline{f_{k}}(v^{-1})$
@@
$[w^{0}]a\rho(a)=\sum q^{-\frac{k^{2}m}{4}}f_{k}(q^{k}v)v^{-\frac{km}{2}}\overline{f_{k}}(v^{-1}).$ (2.9)

### v2

@@ 8 Twisted traces
$q^{k}S\bigg(\frac{-v^{k}}{qv-q^{-1}v^{-1}}F\bigl(v^{-1}\bigr)+\frac{v^{-k}}{q^{-1}v-qv^{-1}}F(v)\bigg)$
$\qquad=q^{-\frac{m}{2}}S\bigg(F(qv)\frac{v^{-k-m}}{v-v^{-1}}-F\bigl(qv^{-1}\bigr)\frac{v^{k+m}}{v-v^{-1}}\bigg).$ (8.1)
Conversely, we claim that every $S$ satisfying (8.1) gives a trace $T$ by $T\bigl(\sum w^{i}f_{i}\bigr)=S(f_{0})$. The algebra $\mathcal{A}$ is generated by $H_{0}$, $H_{1}$ and $v+v^{-1}$. Hence, it is enough check the trace condition T(ba)=T(ag(b)){T(ba)=T(ag(b))} for $b=H_{0},H_{1}$ or $b=v+v^{-1}$. In the case when $b=v+v^{-1}$, the trace condition is satisfied for any $S$. In the case when $b=H_{k}$, it is enough to take $a\in\mathcal{A}_{1}$. In this case, the computation above shows that $T(ba)=T(ag(b))$ is equivalent to (8.1).
Now we check that the integral from the statement of the proposition satisfies the equation (8.1)
$q^{k}\left(\int_{\varepsilon S^{1}}\left(\frac{-z^{k}}{qz-q^{-1}z^{-1}}F\bigl(z^{-1}\bigr)+\frac{z^{-k}}{q^{-1}z-qz^{-1}}F(z)\right)\omega(z)\frac{{\rm d}z}{z}\right)$
$\qquad=q^{-\frac{m}{2}}\left(\int_{\varepsilon S^{1}}\left(F(qz)\frac{z^{-k-m}}{z-z^{-1}}-F\bigl(qz^{-1}\bigr)\frac{z^{k+m}}{z-z^{-1}}\right)\omega(z)\frac{{\rm d}z}{z}\right).$
@@
$\int_{\varepsilon S^{1}}\bigg(F(qz)\frac{z^{-k-m}}{z-z^{-1}}-F\bigl(qz^{-1}\bigr)\frac{z^{k+m}}{z-z^{-1}}\bigg)\omega(z)\frac{{\rm d}z}{z}$
$\qquad=\int\limits_{q^{-1}\varepsilon S^{1}}F(qz)\frac{z^{-k-m}}{z-z^{-1}}\omega(z)\frac{{\rm d}z}{z}-\int\limits_{q\varepsilon S^{1}}F\bigl(qz^{-1}\bigr)\frac{z^{k+m}}{z-z^{-1}}\omega(z)\frac{{\rm d}z}{z}$
$\qquad=\int_{\varepsilon S^{1}}F(z)\frac{q^{k+m}z^{-k-m}}{q^{-1}z-qz^{-1}}\omega(q^{-1}z)\frac{{\rm d}z}{z}-\int_{\varepsilon S^{1}}F\bigl(z^{-1}\bigr)\frac{q^{k+m}z^{k+m}}{qz-q^{-1}z^{-1}}\omega(qz)\frac{{\rm d}z}{z}$
$\qquad=q^{k+\frac{m}{2}}\bigg(\int_{\varepsilon S^{1}}F(z)\frac{z^{-k}}{q^{-1}z-qz^{-1}}\omega(z)\frac{{\rm d}z}{z}-\int_{\varepsilon S^{1}}F\bigl(z^{-1}\bigr)\frac{z^{k}}{qz-q^{-1}z^{-1}}\omega(z)\frac{{\rm d}z}{z}\bigg).$
@@
This shows that every function $\omega$ satisfying
$\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z),\qquad\omega(z)=\omega\bigl(z^{-1}\bigr)$
gives a trace. The space of such functions has dimension $1+\frac{m}{2}$. To show that, consider $w(z)=\omega\bigl({\rm e}^{2\pi{\rm i}z}\bigr)$. Let $q={\rm e}^{2\pi{\rm i}\tau}$. Then
$w(z+1)=w(z),\qquad w(-z)=w(z)\qquad\text{and}\qquad w(z+\tau)={\rm e}^{-\pi\mathrm{i}m(\tau+2z)}w(z).$
Any such function can be obtained as $C\prod_{j=1}^{\frac{m}{2}}\vartheta(z-\alpha_{j})\vartheta(z+\alpha_{j})$, where $\vartheta$ is Jacobi theta function and $\alpha_{j}$ and $C$ are any numbers.
Now we prove that all maps $S$ satisfying (8.1) can be obtained in this way. It is enough to prove that the dimension of the space of possible $S$ is at most $\frac{m}{2}+1$. Since any $S$ satisfying (8.1) defines a trace $T$, we will use the trace condition below.
Taking regular $F$ such that $F(q)=0$ and $F(-q)\neq 0$, we see that the argument of $S$ on the left-hand side of (8.1) has a pole at $v=-q^{\pm 1}$ and is regular at $v=q^{\pm 1}$. The argument on the right-hand side of (8.1) is regular. Hence, $S\bigl(\frac{1}{(v+q)(v^{-1}+q)}\bigr)$ is defined by the restriction of $S$ to $\mathbb{C}\big[v+v^{-1}\big]$. Taking regular $F$ such that F(q)=0F(-q)=0 and $F(q)\neq 0$, we get that $S\bigl(\frac{1}{(v-q)(v-q^{-1})}\bigr)$ is defined by the restriction of $S$ to $\mathbb{C}\big[v+v^{-1}\big]$. Now, suppose that $F$ has a pole at $q^{k}$ for some $k\neq 1$. Then the argument of $S$ on the left-hand side of (8.1) has a pole at $\pm q^{\pm 1}$ and $q^{\pm k}$ (no pole if k=0{k=0}) and the right-hand side has a pole at $q^{\pm(k-1)}$. It follows by induction with the base case $k=1$ that $S\bigl(\frac{1}{(v-q^{k})(v^{-1}-q^{k})}\bigr)$ is defined by the restriction of $S$ to $\mathbb{C}\big[v+v^{-1}\big]$. A similar argument shows that $S\bigl(\frac{1}{(v+q^{k})(v^{-1}+q^{k})}\bigr)$ is defined by the restriction of $S$ to $\mathbb{C}\big[v+v^{-1}\big]$.
Take the same $a\in\mathcal{A}$ as before and $b=-w+w^{-1}$. Writing $T(ba)=T(ag(b))$ in terms of $S$, we get
$S\bigl(F(v)+F\bigl(v^{-1}\bigr)\bigr)=q^{-\tfrac{m}{2}}S\bigl(F(qv)v^{-m}+F\bigl(qv^{-1}\bigr)v^{m}\bigr).$
For $F(v)=v^{k}$, we get
$S\bigl(v^{k}+v^{-k}\bigr)=q^{k-\tfrac{m}{2}}S\bigl(v^{k-m}+v^{m-k}\bigr).$
It follows that $S$ is uniquely defined by $S(1),S\bigl(v+v^{-1}\bigr),\dots,S\bigl(v^{\frac{m}{2}}+v^{-\frac{m}{2}}\bigr)$. Therefore, the dimension of the space of linear maps $S$ satisfying (8.1) is at most $1+\frac{m}{2}$, as claimed. ∎
9 Positive traces
Proposition 9.1.
Let $T$ be a positive trace. Let $\omega$ be a function from Proposition 8.2. Then $\omega(1)=\omega(-1)=0$.
@@ 9 Positive traces
$S(F(z))=\int_{S^{1}}F(z)\omega(z)\frac{{\rm d}z}{z}+\sum_{i>0}c_{i,\pm}\operatorname{Res}_{\pm q^{i}}F,$
where $c_{i,\pm}$ are nonzero multiples of $\omega(\pm 1)$, so that for at least one of the choices of sign all $c_{i,\pm}$ are nonzero. Without loss of generality, assume that $c_{i}=c_{i,+}$ are nonzero.
Let $a=\bigl(w^{2}F(v)-w^{-2}vF\bigl(v^{-1}\bigr)\bigr)(v-1)^{-1}$, where $F(v)$ is regular. This is an element of $\mathcal{A}$. Then
$\rho(a)=q^{m}\bigl(-w^{2}v^{1-m}\overline{F}(v)+w^{-2}v^{m}\overline{F}\bigl(v^{-1}\bigr)\bigr)(v-1)^{-1}$
@@
$\big[w^{0}\big]a\rho(a)={}$ $q^{m}\bigl(w^{2}F(v)(v-1)^{-1}w^{-2}v^{m}\overline{F}\bigl(v^{-1}\bigr)(v-1)^{-1}$
$+w^{-2}vF\bigl(v^{-1}\bigr)(v-1)^{-1}w^{2}\overline{F}(v)v^{1-m}(v-1)^{-1}\bigr).$
@@
$w^{2}F(v)(v-1)^{-1}w^{-2}v^{m}\overline{F}\bigl(v^{-1}\bigr)(v-1)^{-1}+w^{-2}vF\bigl(v^{-1}\bigr)(v-1)^{-1}w^{2}v^{1-m}\overline{F}(v)(v-1)^{-1}$
$\qquad=F\bigl(q^{2}v\bigr)\overline{F}\bigl(v^{-1}\bigr)\bigl(q^{2}v-1\bigr)^{-1}v^{m}(v-1)^{-1}$
$\phantom{\qquad=}{}+F\bigl(q^{2}v^{-1}\bigr)\overline{F}(v)q^{-2}v\bigl(q^{-2}v-1\bigr)^{-1}v^{1-m}(v-1)^{-1}=G(v)+G\bigl(v^{-1}\bigr),$
@@
$G(v)=\overline{F}\bigl(v^{-1}\bigr)F\bigl(q^{2}v\bigr)\cdot\frac{v^{m}}{\bigl(q^{2}v-1\bigr)(v-1)}$
@@
$G\bigl(v^{-1}\bigr)=\overline{F}(v)F\bigl(q^{2}v^{-1}\bigr)\cdot\frac{q^{-2}v^{2-m}}{(q^{-2}v-1)(v-1)}.$
@@
$T(a\rho(a))=S\bigl(\big[w^{0}\big]a\rho(a)\bigr)=S\bigl(G(v)+G\bigl(v^{-1}\bigr)\bigr)=\int_{\varepsilon S^{1}}\bigl(G(z)+G\bigl(z^{-1}\bigr)\bigr)\omega(z)\frac{{\rm d}z}{z}.$ (9.1)
@@
$\overline{F}(1)\frac{F\bigl(q^{2}\bigr)}{q^{2}-1}.$
The only pole of $G\bigl(z^{-1}\bigr)$ between 0 and $qS^{1}$ is $v=q^{2}$, and the residue is
$\overline{F}(q^{2})\cdot\frac{F(1)q^{-2}q^{2(2-m)}}{\bigl(q^{2}-1\bigr)q^{-2}}=F(1)\overline{F}(q^{2})\frac{q^{2(2-m)}}{q^{2}-1}.$
Therefore,
$\int_{\varepsilon S^{1}}\bigl(G(z)+G\bigl(z^{-1}\bigr)\bigr)\omega(z)\frac{{\rm d}z}{z}=$ $c_{0}\overline{F}(1)\frac{F\bigl(q^{2}\bigr)}{q^{2}-1}+c_{1}F(1)\overline{F}\bigl(q^{2}\bigr)\frac{q^{2(2-m)}}{q^{2}-1}$
$+\int_{q^{-1}S^{1}}G(z)\omega(z)\frac{{\rm d}z}{z}+\int_{qS^{1}}G\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}.$ (9.2)
We rewrite the integrals
$\int_{q^{-1}S^{1}}G(z)\omega(z)\frac{{\rm d}z}{z}$ $=\int_{S^{1}}G\bigl(q^{-1}z\bigr)\omega\bigl(q^{-1}z\bigr)\frac{{\rm d}z}{z}$
$=\int_{S^{1}}\overline{F}\bigl(qz^{-1}\bigr)F(qz)\frac{q^{-m}v^{m}}{(qv-1)\bigl(q^{-1}v-1\bigr)}\omega\bigl(q^{-1}z\bigr)\frac{{\rm d}z}{z},$
$\int_{qS^{1}}G\bigl(z^{-1}\bigr)\omega(z)\frac{{\rm d}z}{z}$ $=\int_{S^{1}}G\bigl(q^{-1}z^{-1}\bigr)\omega(qz)\frac{{\rm d}z}{z}$
$=\int_{S^{1}}\overline{F}(qz)F\bigl(qz^{-1}\bigr)\cdot\frac{q^{-m}v^{z-m}}{(v-1)\bigl(q^{2}v-1\bigr)}\cdot\omega(qz)\frac{{\rm d}z}{z}.$
Now take $F(z)=q^{-N}z^{N}-a$ for some constant $a$ and natural number $N$. For large $N$, the leading term of
$c_{0}\overline{F}(1)\frac{F\bigl(q^{2}\bigr)}{q^{2}-1}+c_{1}F(1)\overline{F}(q^{2})\frac{q^{2(2-m)}}{q^{2}-1}$
is $C_{0}aq^{-N}+C_{1}\overline{a}q^{-N}$, where $C_{0}$, $C_{1}$ are constant numbers. For some choice of $a$, this is a negative number with absolute value growing as $q^{-N}$.
On the other hand, for $z\in S^{1}$ polynomials $\overline{F}\bigl(qz^{-1}\bigr)$, $F(qz)$, $\overline{F}(qz)$, $F\bigl(qz^{-1}\bigr)$ have absolute value at most $1+\lvert a\rvert$. Hence, the integrals in the right-hand side of (9.2) are bounded when $N$ tends to infinity. On the other hand, the first two terms in the right-hand side of (9.2) tend to minus infinity. Comparing (9.1) and (9.2), we see that for large enough $N$ we have $T(a\rho(a))<0$, a contradiction with positivity. ∎
In particular, for $m=4$ we get that the function $\omega$ is unique up to a constant. Indeed, consider the function $w$ such that $w(z)=\omega\bigl({\rm e}^{2\pi{\rm i}z}\bigr)$. Then, using $w(0)=w(\tfrac{1}{2})=0$, we get $w(z)=\vartheta\bigl(z-\frac{\tau}{2}\bigr)\vartheta\bigl(z+\frac{\tau}{2}\bigr)\vartheta\bigl(z-\tfrac{1}{2}-\frac{\tau}{2}\bigr)\vartheta\bigl(z+\frac{1}{2}+\frac{\tau}{2}\bigr)$ up to a constant. Here $\vartheta(z)$ is the Jacobi theta function.
Proposition 9.2.
Let $T$ be a positive trace. Then $T$ vanishes on $\mathcal{A}_{1}$.
@@
Using Proposition 8.1, it is enough to prove that $T(a)=0$ for elements $a$ of the form $a=wF(v)-wF\bigl(v^{-1}\bigr)$. For these elements, we have $T(a)=C_{+}F(\sqrt{q})+C_{-}F(-\sqrt{q})$. Assume that $C_{+}$ is nonzero.
Let $a=wF(v)-w^{-1}F\bigl(v^{-1}\bigr)$ for some Laurent polynomial $F$, $b=a+1$. Then
$\rho(b)=q^{-\frac{m}{4}}\bigl(-wv^{\frac{m}{2}}\overline{F}(v)+w^{-1}v^{-\frac{m}{2}}\overline{F}\bigl(v^{-1}\bigr)\bigr)+1.$
Note that $T(b\rho(b))$ depends only on the $w^{0}$ and $w^{1}$ coefficients of $b\rho(b)$. They can be computed as follows
$\big[w^{0}\big]b\rho(b)=q^{-\frac{m}{4}}\bigl(F(qv)v^{-\frac{m}{2}}\overline{F}\bigl(v^{-1}\bigr)+F\bigl(qv^{-1}\bigr)v^{\frac{m}{2}}\overline{F}(v)\bigr)+1,$
$\big[w^{1}\big]b\rho(b)=F(v)-v^{\frac{m}{2}}\overline{F}(v).$
Let $G=F(\sqrt{q}v)$, $I(G)=\int_{S^{1}}\lvert G(z)\rvert^{2}\frac{{\rm d}z}{z}$. We have $S\bigl(\big[w^{0}\big]b\rho(b)\bigr)\leq cI(G)$ for some $c>0$. On the other hand, trace for the odd powers of $w$ is a linear combination of $G(1)$, G(1)G(-1), $\overline{G}(1)$ and $\overline{G}(-1)$; and $G(1)$ has nonzero coefficient. Let $G_{N}(z)=a\bigl(\frac{z+1}{2}\bigr)^{N}$. Then $G_{N}(1)=a$, $G_{N}(-1)=0$. On the other hand, $G_{N}(z)$ is a sequence of functions such that $\lvert G_{N}(z)\rvert\leq 1$ tends to zero for almost all $z$ as $N$ tends to infinity. Since pointwise convergence implies convergence in measure for spaces of finite total measure and the functions $G_{N}$ are uniformly bounded, we deduce that $I(G_{N}(z))$ tends to zero as $N$ tends to infinity. We get that $T(b\rho(b))$ tends to a linear combination of $a$ and $\overline{a}$. We can choose $a$ so that this linear combination does not belong to $\mathbb{R}_{\geq 0}$. We get a contradiction with positivity of $T$. ∎
Let $T$ be a positive trace. Combining Propositions 9.1 and 9.2, it follows that for $a=\sum w^{i}f_{i}$ we have $T(a)=\int f_{0}(z)\omega(z)\frac{{\rm d}z}{z}$ and $\omega(1)=\omega(-1)=1$.
Proposition 9.3.
Let $T$ be a trace on $\mathcal{A}$ of the form $T(a)=\int f_{0}(z)\omega(z)\frac{{\rm d}z}{z}$, where $w$ is a holomorphic function on $\mathbb{C}^{*}$ such that $\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z)$, $\omega(z)=\omega\bigl(z^{-1}\bigr)$ and $\omega(1)=\omega(-1)=0$. Then $T$ is positive if and only if $\omega(z)$ and $z^{-\frac{m}{2}}\omega\bigl(\sqrt{q^{-1}}z\bigr)$ are nonnegative on $S^{1}$.
@@
$\rho(a)=\sum q^{-\frac{k^{2}m}{4}}w^{-k}v^{-\frac{km}{2}}\overline{f_{k}}\bigl(v^{-1}\bigr)$
@@
$\big[w^{0}\big]a\rho(a)=\sum q^{-\frac{k^{2}m}{4}}f_{k}(q^{k}v)v^{-\frac{km}{2}}\overline{f_{k}}\bigl(v^{-1}\bigr).$ (9.3)